Compute the electric force between two charges of 5×10-9 C and -3×10-8 C which are separated by r=10 cm


1
Expert's answer
2022-03-02T14:39:21-0500

The force is given as follows:


"F = k\dfrac{|q_1q_2|}{r^2}"

where "k = 9\times 10^9N\cdot m^2/C^2, q_1 = 5\times 10^{-9}C,q_2 = -3\times 10^{-8}C" and "r = 10cm = 0.1m". Thus, obtain:



"F = \dfrac{9\times10^9\cdot 5\times 10^{-9}\cdot 3\times 10^{-8}}{0.1^2} =1.35\times 10^{-4}N"

Answer. "1.35\times 10^{-4}N".